# Monte Carlo sampling with integrator snippets

### v1

@@ 2.2 Justification outline
Consequently
is an unbiased estimator of $Z_{P}/Z_{0}$.
Lemma 4.
  1. 1.
    $\bar{\mu}_{n-1}\bar{M}_{n}=\mu_{n-1}$,
  2. 2.
    there exists $\bar{L}_{n-1}\colon\mathsf{Z}\times\mathscr{Z}\rightarrow[0,1]$ in (13) such that for any $A,B\in\mathscr{Z}$
  3. 3.
    the importance weight in (14) is well defined and
Proof.
For $(z,A)\in\mathsf{Z}\times\mathscr{Z}$, we let
and note that almost surely $\bar{L}_{n-1}(z,A)\in[0,1]$ and $\bar{L}_{n-1}(z,\mathsf{Z})=1$. For the third statement note that from the second statement, for $A,B\in\mathscr{Z}$ we have $\bar{\mu}_{n-1}\otimes\bar{M}_{n}\big{(}A\times B\big{)}=\bar{\mu}_{n-1}\bar{M}_{n}\accentset{\curvearrowleft}{\otimes}\bar{L}_{n-1}\big{(}A\times B\big{)}$ and from Fubini’s and the $\pi-\lambda$ theorem [7, Theorems 3.1 and 3.2] $\bar{\mu}_{n-1}\otimes\bar{M}_{n}$ and $\bar{\mu}_{n-1}\bar{M}_{n}\accentset{\curvearrowleft}{\otimes}\bar{L}_{n-1}$ are probability distributions coinciding on $\mathscr{Z}\otimes\mathscr{Z}$. To conclude proof of the third statement, for $f\colon\mathsf{Z}^{2}\rightarrow[0,1]$ measurable, we successively apply the definition of the Radon-Nikodym derivative, Fubini’s theorem, use that $\mu_{n-1}\gg\bar{\mu}_{n}$, the first statement of this lemma, Fubini again, the second statement
2.3 Direct extensions
It should be clear that the algorithm we have described lends itself to numerous generalizations, which we briefly discuss below.
There is no reason to limit the number of snippets arising from a seed particle to one, therefore offering the possibility to take advantage of parallel machines. For example the velocity of a given seed particle can be refreshed multiple times, resulting in partial copies of the seed particle from which integrator snippets can be grown.
The main scenario motivating this work, corresponds to the choice, for $n\in\llbracket 0,P\rrbracket$, of$\{\psi_{n,k}=\psi_{n}^{k},k\in\llbracket 0,T\rrbracket\}$ for a given $\psi_{n}\colon\mathsf{Z}\rightarrow\mathsf{Z}$. As should be apparent from the theoretical justification this can be replaced by a general family of invertible mappings $\{\psi_{n,k}\colon\mathsf{Z}\rightarrow\mathsf{Z},k\in\llbracket 0,T\rrbracket\}$ where the $\psi_{n,k}$’s are now not required to be measure preserving in general, in which case the expression for $w_{n,k}(z)$ may involve additional terms of the “Jacobian” type. These mappings may correspond to integrators other than those of Hamilton’s equations but may be more general deterministic mappings and $T$ may have no temporal meaning and only represent the number of deterministic transformations used in the algorithm. More specifically, asssuming that $\upsilon\gg\mu_{n}$ and $\upsilon\gg\upsilon^{\psi_{n,k}^{-1}}$ for some $\sigma$-finite dominating measure $\upsilon$ and letting $\mu_{n}(z):={\rm d}\mu_{n}/{\rm d}\upsilon(z)$ the required weights are now of the form (see Lemmas 32-34)
which still possesses the fundamental property that if $z\sim\bar{\mu}$ (resp. $(i,z)\sim\bar{\mu}$), then with $(i,k)\sim\bar{\nu}(k,i\mid z)$ (resp. $k\sim\bar{\nu}(k\mid i,z)$) we have $\psi_{i,k}(z)\sim\mu$. It is possible to aim to sample from $\bar{\mu}({\rm d}z)$, in which case the pair $(i,k)$ plays the rôle played by $k$ in the earlier simpler scenario. However, in order to introduce the sought persistency, that is use either $\psi_{1}$ or $\psi_{2}$ when constructing a snippet, we focus on the scenario where the pair $(i,z)$ plays the rôle played by $z$ earlier. In other words the target distribution is now $\bar{\mu}(i,{\rm d}z)$, which is to $\mu(i,{\rm d}z):=\gamma_{i}\cdot\mu({\rm d}z)$ what $\bar{\mu}_{n}({\rm d}z)$ in (24) was to $\mu_{n}({\rm d}z)$. The mutation kernel corresponding to (24) is given by
with now the requirement that $\bar{\nu}R(i,{\rm d}z)=\nu(i,{\rm d}z)=\gamma_{i}\cdot\nu({\rm d}z)$. A sensible choice seems to be $R(i,z;j,{\rm d}z^{\prime})=\gamma_{j}\cdot R_{0}(z,{\rm d}z^{\prime})$ with $\nu R_{0}({\rm d}z^{\prime})=\nu({\rm d}z^{\prime})$. With these choices using the near-optimal backward kernel simple substitutions $(i,z)\leftarrow z$ and $(j,z^{\prime})\leftarrow z^{\prime}$ yields
and justifies the earlier abstract presentation.
with $\omega_{n,k}\colon\mathsf{Z}\rightarrow\mathbb{R}_{+}$ for $k\in\llbracket 0,T\rrbracket$ and such that $\sum_{k=0}^{T}\omega_{n,k}(z)=1$ for any $z\in\mathsf{Z}$; this should be contrasted with (8). Such choices ensure that the central identity (11) can be generalized as follows
$\sum_{k=0}^{T}\int f\circ\psi_{n}^{k}(z)\frac{\omega_{n,k}\circ\psi_{n,k}(z)\cdot w_{n,k}(z)}{\sum_{l=0}^{T}\omega_{n,l}\circ\psi_{n,l}(z)w_{n,l}(z)}\bar{\mu}_{n}({\rm d}z)$ $=\sum_{k=0}^{T}\int f\circ\psi_{n,k}(z)\,\bar{\mu}_{n}(k\mid z)\bar{\mu}_{n}({\rm d}z)$
$=\int f(z)\mu_{n}({\rm d}z)\,.$
Note the analogy with the identity behind umbrella sampling [37, 39]. In the light of the developments of Subsection 4.3.2, a potentially useful choice could be $\omega_{n,k}(z)\propto\|z-\psi_{n,k}^{-1}(z)\|$, which however requires additional computations as
$\omega_{n,k}\circ\psi_{n,k}(z)=\frac{\|\psi_{n,k}(z)-z\|}{\sum_{l=0}^{T}\|\psi_{n,k}(z)-\psi_{n,k}\circ\psi_{n,l}^{-1}(z)\|}$
which may require computation of additional states for $l>k$.
We leave exploration of some of these generalisations for future work, although we think that the choice of the leapfrog integrator is particularly natural and attractive here.
2.4 Rational and computational considerations
The benefit of using all states along integrator snippets in sampling algorithms has been noted in the literature. For example, in the context of Hamiltonian integrators, with $H_{n}(z):=-\log\mu_{n}(z)$ and $\psi_{n,k}=\psi_{n}^{k}$, it is known that for $z\in\mathsf{Z}$ the mapping $k\mapsto H_{n}\circ\psi_{n}^{k}(z)$ is typically oscillatory, motivating for example the x-tra chance algorithm of [11]. The “windows of state” strategy of [28, 29] effectively makes use of the mixture $\bar{\mu}$ as an instrumental distribution and improved performance is noted, with performance seemingly improving with dimension on particular problems; see Section 5 for a more detailed discussion. Further averaging is well known to address scenarios where components of $x_{t}$ evolve on different scales and no choice of a unique integration time $\tau:=T\times\varepsilon$ can accommodate all scales [29, 22, section 3.2]; averaging addresses this issue effectively, see Example 27. We also note that, keeping $T\times\varepsilon$ constant, in the limit as $\varepsilon\rightarrow 0$ the average in (14) corresponds to the numerical integration, Riemann like, along the contour of $H_{n}(z)$, effectively leading to some form of Rao-Blackwellization of this contour; this is discussed in detail in Section 4.
Another benefit we have observed with the SMC context is the following. Use of a WF-SMC strategy involves comparing particles within each Markov snippet arising from a single seed particle, while our strategy involves comparing all particles across snippets, which proves highly beneficial in practice. This seems to bring particular robustness to the choice of the integrator parameters $\varepsilon$ and $T$ and can be combined with highly robust adaptation schemes taking advantage of the population of samples, in the spirit of [21, 22]; see Subsections 2.5 and 2.6.
2.5 Numerical illustration: logistic regression
$L(x)=\prod_{i}^{n}F(z_{i}^{\top}x\cdot y_{i}),$ (17)
For both algorithms, we set the total number of particles at each SMC step to be $N(T+1)=10,000$. For the waste-free SMC, the Markov kernel is chosen to be a random-walk Metropolis-Hastings kernel with covariances adaptively computed as $2.38^{2}/d\ \hat{\Sigma}$, where $\hat{\Sigma}$ is the empirical covariance matrix obtained from the particles in the previous SMC step. For the Hamiltonian Snippet-SMC algorithm, we set $\psi_{n}$ to be the one-step leap-frog integrator with stepsize $\varepsilon$, $U_{n}(x)=-\log(\pi_{n}(x))$ and $\varpi$ a $\mathcal{N}(0,\mathrm{Id})$. To investigate the stability of our algorithm, we ran Hamiltonian Snippet SMC with $\varepsilon=0.05,0.1,0.2$ and $0.3$. For both algorithms, the temperatures $\lambda_{n}$ are adaptively chosen so that the effective sample size (ESS) of the current SMC step will be $\alpha ESS_{max}$, where $ESS_{max}$ is the maximum ESS achievable at the current step. In our experiments, we have chosen $\alpha=0.3,0.5$ and $0.7$ for both algorithms.
Performance comparison
Figure 2 shows the boxplots of estimates of the logarithm of the normalising constant obtained from both algorithms, for different choices of $N$ and $\varepsilon$ for the Hamiltonian Snippet SMC algorithm. The boxplots are obtained by running both algorithms 100 times for different of the algorithm parameters, with $\alpha=0.5$ in all setups. Several points are worth observing. For a suitably choice of $\varepsilon$, the Hamiltonian Snippet SMC algorithm can produce stable and consistent estimates of the normalising constant with $10,000$ particles at each iteration. On the other hand, however, the waste-free SMC algorithm fails to produce accurate results for the same computational budget. It is also clear that with larger values of $N$ (meaning smaller value of $T$ and hence shorter snippets), the waste-free SMC algorithm produces results with larger biases and variability. For Hamiltonian Snippet SMC algorithm, the results are stable both for short and long snippets when $\varepsilon$ is equal to $0.1$ or $0.2$. Another point is that when $\varepsilon=0.05$ or $0.3$, the Hamiltonian Snippet SMC algorithm becomes unstable with short (i.e. $\varepsilon=0.05$) and long (i.e. $\varepsilon=0.3$). Possible reasons are for too small a stepsize the algorithm is not able to explore the target distribution efficiently, resulting in unstable performances. On the other hand, when the stepsize is too large, the leapfrog integrator becomes unstable, therefore affecting the variability of the weights; this is a common problem of HMC algorithms. Hence, a long trajectory will result in detoriate estimations. Hence, to obtain the best performance, one should find a way of tuning the stepsize and trajectory length along with the SMC steps.
Figure 2: Estimates of the normalising constant (in log scale) obtained from both Hamiltonian Snippet SMC and waste-free SMC algorithm. Left: Estimate obtained from Hamiltonian Snippet SMC algorithm with different values of $\varepsilon$. Right: Estimates obtained from waste-free SMC algorithm.
Figure 3: Estimates of the mean of marginals obtained from both Hamiltonian Snippet SMC and the waste-free SMC algorithms. Left: Estimate obtained from the Hamiltonian Snippet SMC algorithm with difference values of $\varepsilon$. Right: Estimates obtained from the waste-free SMC algorithm.
Computational Cost
Figure 4: Simulation times for both algorithms. Left: Simulation time for Hamiltonian Snippet SMC algorithm, with $N=100,\varepsilon=0.2,\alpha=0.5$. Right:Simulation times for the waste-free SMC with $N=100$
2.6 Numerical illustration: simulating from filamentary distributions
:=$\mathcal{M}:=\left\{x\in\mathsf{X}:\ell(x)=0\right\},$
$\pi_{\epsilon}(x)\propto k_{\epsilon}\circ\ell(x)\,\pi(x)\,.$
[…]
Integrator Snippet SMC
[…]
2.7 Links to the literature
Readers familiar with the “waste-free SMC” algorithm (WF-SMC) of [14] where, similarly to Alg. 2, seed particles are extended with an MCMC kernel leaving $\mu_{n-1}$ invariant at iteration $n$, yielding $N\times(T+1)$ particles subsequently whittled down to $N$ new seed particles. A first difference is that while generation of an integrator snippet can be interpreted as applying a sequence of deterministic Markov kernels (see Subsection 3.3 for a detailed discussion) what we show is that the mutation kernels involved do not have to leave $\mu_{n-1}$ invariant; in fact we show in Section 3 that this indeed is not a requirement, therefore offering more freedom. Further, it is instructive to compare our procedure with the following two implementations of WF-SMC using an HMC kernel (4) for the mutation. A first possibility for the mutation stage is to run $T$ steps of an HMC update in sequence, where each update uses one integrator step. Assuming no velocity refreshment along the trajectory this would lead to the exploration of a random number of states of our integrator snippet due to the accept/reject mechanism involved; incorporating refreshment or partial refreshment would similarly lead to a random number of useful samples. Alternatively one could consider a mutation mechanism consisting of an HMC build around $T$ integration steps and where the endpoint of the trajectory would be the mutated particle; this would obviously mean discarding $T-1$ potentially useful candidates. To avoid possible reader confusion, we note apparent typos in [14, Proposition 1], provided without proof, where the intermediate target distribution of the SMC algorithms, $\bar{\mu}_{n}$ in our notation (see (18)), seems improperly stated. The statement is however not used further in the paper.
$\bar{\mu}({\rm d}z)=\sum_{k\in\mathbb{Z}}\omega_{k}\mu_{k}({\rm d}z)\,.$
Much closer in spirit to our work is the “window of states” idea proposed in [28, 29] in the context of HMC algorithms. Although the MCMC algorithms of [28, 29] ultimately target $\mu$ and are not reversible, they involve a reversible MH update with respect to $\bar{\mu}$ and additional transitions permitting transitions from $\mu$ to $\bar{\mu}$ and $\bar{\mu}$ to $\mu$; see Section 5 for full details.
A link we have not explored in the present manuscript is that to normalizing flows [36, 26] and related literature. In particular the ability to tune the parameter of the leapfrog integrator suggests that our methodology lends itself learning normalising flows. We finally note an analogy of such approaches with umbrella sampling [37, 39], although the targetted mixture is not of the same form, and the “Warp Bridge Sampling” idea of [27].
3 Sampling Markov snippets
@@
In this section we develop the Markov snippet framework, largely inspired by the WF-SMC framework of [14] but provide here a detailed derivation following the standard SMC sampler framework [16] which allows us to consider much more general mutation kernels; integrator snippet SMC is recovered as a particular case. Importantly we provide recipes to compute some of the quantities involved using simple criteria (see Lemma 9 and Corollary 10) which allow us to consider unusual scenarios such as in Subsection 3.4.
3.1 Markov snippet SMC sampler or waste free SMC with a difference
@@
Note that [14] set $R_{n}=M_{n}$, which we do not want in our later application and further assume that $M_{n}$ is $\mu_{n-1}-$invariant, which is not necessary and too constraining for our application in Section 3.3 (corresponding to the application in the introductory Subsection 2.2). We only require the condition $\mu_{n-1}R_{n}=\mu_{n-1}$ is required. As in Subsection 2.2 we consider the optimal backward kernel $\bar{L}_{n}\colon\mathsf{Z}^{T+1}\times\mathscr{Z}^{\otimes(T+1)}\rightarrow[0,1]$, given for $(\mathsf{z},A)\in\mathsf{Z}^{T+1}\times\mathscr{Z}^{\otimes(T+1)}$ by
@@
The corresponding folded version of the algorithm is given in Alg. 4, with $\check{\mathsf{z}}_{n}:=(\check{z}_{n,0},\check{z}_{n,1},\ldots,\check{z}_{n,T})$…:
@@
3.2 Theoretical justification
In this section we provide the theoretical justification for the correctness of Alg. 4 and an alternative proof for Alg. 3, seen as a particular case of Alg. 4.
@@ 3.2 Theoretical justification
The following result is important in two respects. First it establishes that if $M,L$ satisfy a simple property then $w_{k}$ always have a simple expresssion in terms of certain densities of $\mu$ and $\nu$ – this implies in particular that in Subsection 3.1 the kernel $M_{n}$ is not required to leave $\mu_{n-1}$ invariant to make the method implementable [14]. Second it provides a direct justification of the validity of advanced schemes – see Example 14. This therefore establishes that generic and widely applicable sufficient conditions for $\bar{w}(\mathsf{z},\mathsf{z}^{\prime})$ to be tractable are $\nu R=\nu$ and the notion of $(\upsilon,M,M^{*})$-reversibility.
@@
where we have used Lemma  34.
@@
3.3 Revisiting sampling with integrator snippets
In this scenario we have $\mu({\rm d}z)=\pi({\rm d}x)\varpi({\rm d}v)$ assumed to have a density w.r.t. a $\sigma$-finite measure $\upsilon$, and $\psi$ is an invertible mapping $\psi\colon\mathsf{Z}\rightarrow\mathsf{Z}$ such that $\upsilon^{\psi}=\upsilon$ and $\psi^{-1}=\sigma\circ\psi\circ\sigma$ with $\sigma\colon\mathsf{Z}\rightarrow\mathsf{Z}$ such that $\mu\circ\sigma(z)=\mu(z)$. In his manuscript we focus primarily on the scenario where $\psi$ is a discretization of Hamilton’s equations for a potential $U\colon\mathsf{X}\rightarrow\mathbb{R}$ e.g. a leapfrog integrator. We consider now the scenario where, in the framework developed in Section 3.1, we let $M(z,{\rm d}z^{\prime}):=\delta_{\psi(z)}\big{(}{\rm d}z^{\prime}\big{)}$ be the deterministic kernel which maps the current state $z\in\mathsf{Z}$ to $\psi(z)$. Define $\Psi(z,{\rm d}z^{\prime}):=\delta_{\psi(z)}({\rm d}z^{\prime})$ and $\Psi^{*}(z,{\rm d}z^{\prime}):=\delta_{\psi^{-1}(z)}({\rm d}z^{\prime})$; we exploit the ideas of [1, Proposition 4] to establish that $\Psi^{*}$ is the $\upsilon-$adjoint of $\Psi$ if $\upsilon$ is invariant under $\psi$.
@@ 3.3 Revisiting sampling with integrator snippets
hence recovering the expression used in Section 2. This together with the results of Section 3.1 provides an alternative justification of correctness of Alg. 3 and hence Alg. 2. Note that this choice for $L_{\mu}$ corresponds to the so-called optimal scenario; this can be deduced from Lemma 11 or by noting that
@@
3.4 More complex integrator snippets
@@
for some unit length vector field $n\colon\mathsf{X}\rightarrow\mathsf{X}$. This can therefore be used as part of Alg. 4; care must be taken when compute the weights $w_{n,k}$ and $\bar{w}_{n}$, see Section 3-3.3 provide the tools to achieve this. For example, in the situation where $\upsilon$ is the Lebesgue measure on $\mathsf{Z}=\mathbb{R}^{d}$ and $\psi_{2}={\rm Id}$ then we recover Alg. 3 or equivalently Alg. 2.
@@
The corresponding kernel $M^{\otimes T}$ is described algorithmically in Alg. 5. In the situation where $C:=\{x\in\mathsf{X}\colon c(x)=0\}$ for a continuously differentiable function $c\colon\mathsf{X}\rightarrow\mathbb{R}$, the bounces described in (25) can be defined in terms of the field $x\mapsto n(x)$such that
@@
This justifies the ideas of [5], where a process of the type given in Alg. 5 is used as a proposal within a MH update, although the possibility of a rejection after the second stage seems to have been overlooked in that reference.

### v2

@@ Appendix C Proofs for Section 2
Appendix D Sampling Markov snippets
@@
In this section we develop the Markov snippet framework, largely inspired by the WF-SMC framework of [26] but provide here a detailed derivation following the standard SMC sampler framework [28] which allows us to consider much more general mutation kernels; integrator snippet SMC is recovered as a particular case. Importantly we provide recipes to compute some of the quantities involved using simple criteria (see Lemma 18 and Corollary 19) which allow us to consider unusual scenarios such as in Appendix D.4.
D.1 Markov snippet SMC sampler or waste free SMC with a difference
@@
Note that [26] set $R_{n}=M_{n}$, which we do not want in our later application and further assume that $M_{n}$ is $\mu_{n-1}-$invariant, which is not necessary and too constraining for our application in Appendix D.3 (corresponding to the application in the introductory Subsection 2.2). We only require the condition $\mu_{n-1}R_{n}=\mu_{n-1}$ is required. As in Subsection 2.2 we consider the optimal backward kernel $\bar{L}_{n}\colon\mathsf{Z}^{T+1}\times\mathscr{Z}^{\otimes(T+1)}\rightarrow[0,1]$, given for $(\mathsf{z},A)\in\mathsf{Z}^{T+1}\times\mathscr{Z}^{\otimes(T+1)}$ by
@@
The corresponding folded version of the algorithm is given in Alg. 6, with $\check{\mathsf{z}}_{n}:=(\check{z}_{n,0},\check{z}_{n,1},\ldots,\check{z}_{n,T})$…:
@@
D.2 Theoretical justification
In this section we provide the theoretical justification for the correctness of Alg. 6 and an alternative proof for Alg. 3, seen as a particular case of Alg. 6.
@@ D.2 Theoretical justification
The following result is important in two respects. First it establishes that if $M,L$ satisfy a simple property then $w_{k}$ always have a simple expresssion in terms of certain densities of $\mu$ and $\nu$ – this implies in particular that in Appendix D.1 the kernel $M_{n}$ is not required to leave $\mu_{n-1}$ invariant to make the method implementable [26]. Second it provides a direct justification of the validity of advanced schemes – see Example 23. This therefore establishes that generic and widely applicable sufficient conditions for $\bar{w}(\mathsf{z},\mathsf{z}^{\prime})$ to be tractable are $\nu R=\nu$ and the notion of $(\upsilon,M,M^{*})$-reversibility.
@@
where we have used Lemma  12.
@@
D.3 Revisiting sampling with integrator snippets
In this scenario we have $\mu({\rm d}z)=\pi({\rm d}x)\varpi({\rm d}v)$ assumed to have a density w.r.t. a $\sigma$-finite measure $\upsilon$, and $\psi$ is an invertible mapping $\psi\colon\mathsf{Z}\rightarrow\mathsf{Z}$ such that $\upsilon^{\psi}=\upsilon$ and $\psi^{-1}=\sigma\circ\psi\circ\sigma$ with $\sigma\colon\mathsf{Z}\rightarrow\mathsf{Z}$ such that $\mu\circ\sigma(z)=\mu(z)$. In his manuscript we focus primarily on the scenario where $\psi$ is a discretization of Hamilton’s equations for a potential $U\colon\mathsf{X}\rightarrow\mathbb{R}$ e.g. a leapfrog integrator. We consider now the scenario where, in the framework developed in Appendix D.1, we let $M(z,{\rm d}z^{\prime}):=\delta_{\psi(z)}\big{(}{\rm d}z^{\prime}\big{)}$ be the deterministic kernel which maps the current state $z\in\mathsf{Z}$ to $\psi(z)$. Define $\Psi(z,{\rm d}z^{\prime}):=\delta_{\psi(z)}({\rm d}z^{\prime})$ and $\Psi^{*}(z,{\rm d}z^{\prime}):=\delta_{\psi^{-1}(z)}({\rm d}z^{\prime})$; we exploit the ideas of [2, Proposition 4] to establish that $\Psi^{*}$ is the $\upsilon-$adjoint of $\Psi$ if $\upsilon$ is invariant under $\psi$.
@@ D.3 Revisiting sampling with integrator snippets
hence recovering the expression used in Section 2. This together with the results of Appendix D.1 provides an alternative justification of correctness of Alg. 3 and hence Alg. 2. Note that this choice for $L_{\mu}$ corresponds to the so-called optimal scenario; this can be deduced from Lemma 20 or by noting that
@@
D.4 More complex integrator snippets
@@
for some unit length vector field $n\colon\mathsf{X}\rightarrow\mathsf{X}$. This can therefore be used as part of Alg. 6; care must be taken when compute the weights $w_{n,k}$ and $\bar{w}_{n}$, see Appendix D-D.3 provide the tools to achieve this. For example, in the situation where $\upsilon$ is the Lebesgue measure on $\mathsf{Z}=\mathbb{R}^{d}$ and $\psi_{2}={\rm Id}$ then we recover Alg. 3 or equivalently Alg. 2.
@@
The corresponding kernel $M^{\otimes T}$ is described algorithmically in Alg. 7. In the situation where $C:=\{x\in\mathsf{X}\colon c(x)=0\}$ for a continuously differentiable function $c\colon\mathsf{X}\rightarrow\mathbb{R}$, the bounces described in (37) can be defined in terms of the field $x\mapsto n(x)$such that
@@
This justifies the ideas of [8], where a process of the type given in Alg. 7 is used as a proposal within a MH update, although the possibility of a rejection after the second stage seems to have been overlooked in that reference.